OpenStudy (anonymous):

Give the cardinality of the set. (If the set has cardinality No, enter ALEPH NULL.) {44, 45, . . . , 189, 190}

OpenStudy (amistre64):

its still gonna be 190-44+1 .... unless there is something im missing

OpenStudy (anonymous):

That's the answer?

OpenStudy (amistre64):

...... is what the answer?

OpenStudy (anonymous):

190-44+1?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

you might wanna actually do the math on it tho ....

OpenStudy (anonymous):

So how did you get that?

OpenStudy (anonymous):

I'm sorry, I'm just having a difficult time understanding this section :/

OpenStudy (amistre64):

let me show you on something more doable ... ok?

OpenStudy (anonymous):

Okay

OpenStudy (amistre64):

suppose we had the set {4,5,6,7,8} , the cardinality of a set just tells us how many terms are in it. how many terms do you count in this set?

OpenStudy (anonymous):

5

OpenStudy (amistre64):

good, now what happens when we try to find the range of this set; lets subtract the first term from the last term: 8-4 = ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

as we can see, 4 is not 5 is it? so we need to add 1 to adjust for it right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

so lets try your set: {44, 45, . . . , 189, 190} do you see now where i get the 190, the 44, and why i would add 1 to it all?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

good :) then the cardinality of your set is; 190 - 44 + 1

OpenStudy (anonymous):

Okayy, my teacher has us do this n+1 thing?

OpenStudy (amistre64):

n just means "some number" usually the last number in the set

OpenStudy (anonymous):

Can we go through another example so I can be sure?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

{16,20,24,...,400,404} 1) subtract 404 from 16 correct?

OpenStudy (amistre64):

this one is trickier, do you see why?

OpenStudy (anonymous):

because you add 4 right?

OpenStudy (amistre64):

correct, how would you adjust for that?

OpenStudy (anonymous):

404-16+4?

OpenStudy (amistre64):

lol ... i wish; lets try to work it out on something smaller.

OpenStudy (anonymous):

Okay :/

OpenStudy (amistre64):

{4,6,8,10,12} ,this adds by 2 right?

OpenStudy (amistre64):

and we know theres 5 terms

OpenStudy (anonymous):

Correct

OpenStudy (amistre64):

12 - 4 = 8 + 2= 10/2 = 5 maybe? but thats just guessing ... lets try it out on something that is by 3s {4,7,10,13,16} 16-4 = 12+3 = 15/3 = 5 ... looks good so far

OpenStudy (amistre64):

7-4 = 3+3 = 6/3 = 2 .... it seems to work to me

OpenStudy (amistre64):

last - first + difference -------------------- = cardinality difference

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

lets try yours now :)

OpenStudy (anonymous):

Okay :)

OpenStudy (amistre64):

{16,20,24,...,400,404} ; added by 4 404 - 16 + 4 ------------ = cardinality then would you agree? 4

OpenStudy (anonymous):

98?

OpenStudy (amistre64):

101-4 = 97 + 1 = 98; yes

OpenStudy (anonymous):

Where did you get 101 from?

OpenStudy (amistre64):

404/4 = 101 16/4 = 4 and 4/4 = 1

OpenStudy (amistre64):

i just divided it all out first then summed it up

OpenStudy (amistre64):

in the end, i did it the same way you did

OpenStudy (amistre64):

i see another method we could have used; but its not as straight forward as this one

OpenStudy (amistre64):

in order to get the set itself; we could have used the formula: A(n) = 16 +4(n-1); where "n" tells us the number of terms

OpenStudy (anonymous):

How did you get the 101?

OpenStudy (amistre64):

\[\frac{404-16+4}{4}=\frac{404}{4}-\frac{16}{4}+\frac{4}{4}=101-4+1\]

OpenStudy (amistre64):

i seperated the sumation into individual parts and reduced to work with smaller numbers

OpenStudy (amistre64):

A(n) = 16 +4(n-1) ; and we know the last term = 404 soo... 404 = 16 +4(n-1) 404 - 16 = 4(n-1) 404 - 16 -------- = n -1 4 404 - 16 -------- + 1 = n 4 hunh ... its the same formula we worked out above ...

OpenStudy (anonymous):

Okay so the cardinality would be 98?

OpenStudy (amistre64):

yes, 98 is correct

OpenStudy (anonymous):

How would I know if a set had cardinality ALEPH NULL?

OpenStudy (anonymous):

Im unsure about the Alpha Null thing

OpenStudy (amistre64):

id have to look up the terminology, but i think it means its either infinite and cant be counted, or its empty and cant be counted

OpenStudy (amistre64):

Aleph-Null bottles of sode on the wall, Aleph-Null bottles of pop, Take one down, and pass it around, Aleph-null bottles of soda on the wall .... " The set theory symbol "Aleph-Null" refers to a set having the same cardinal number as the "small" infinite set of integers."

OpenStudy (amistre64):

in other words, when the set has infinite terms